REU 2007: SCHEDULE FOR WEEKS SEVEN AND EIGHT

BABAI’S AFTERNOON TALKS ARE IN R276 OTHER TALKS ARE IN ECKHART 206

SEVENTH WEEK

Date SESAME Babai Kerr Angeltveit
Mon, July 30 9–12, 1–4 3:00–4:45  5:00–6:00
Tue, July 31 9–12, 1–4 4:30–6:00
Wed, August 1 9–12, 1–4 3:00–4:45  5:00–6:00
Thu, August 2 9–12, 1–4  4:30–6:00
Fri, August 3 9–12, 1–4 3:00–4:45  5:00–6:00

EIGHTH WEEK

Date SESAME Babai Kerr Angeltveit
Mon, August 6 9–12, 1–4 3:00–4:45  5:00–6:00
Tue, August 7 9–12, 1–4  4:30–6:00
Wed, August 8 9–12, 1–4 3:00–4:45  5:00–6:00
Thu, August 9 9–12, 1–4  4:30–6:00
Fri, August 10 9–12, 1–4 3:00–4:45  5:00–6:00

Here is the schedule in summary form:

SEVENTH WEEK:

SESAME: 9:00–12:00 and 1:00–4:00 M-F Babai (Discrete): 3:00–4:45 MWF (SESAME participants: come on in) Kerr (Toric geometry): TTh 4:30–6:00 Angeltveit (Products on spaces): 5:00–6:00 MWF

EIGHTH WEEK:

SESAME: 9:00–12:00 and 1:00–4:00 M-F Babai (Discrete): 3:00–4:45 MWF (SESAME participants: come on in) Kerr (Algebraic geometry): TTh 4:30–6:00 Angeltveit (Products on spaces): 5:00–6:00 MWF

Expected, August 8-10:

WEDNESDAY, 6:30 pm: Student presentations and dinner THURSDAY, 6:30 pm: Student presentations and dinner FRIDAY, 6:30: Student presentations and dinner